If the length of a rectangle is 3 feet longer than the width and the diagonal is 15 feet, then what are the length and the width?

Answer by checkley77(12844)

x^2+(x+3)^2=15^2
x^2+x^2+6x+9=225
2x^2+6x+9-225=0
2x^2+6x+216
2(x^2+3x-108)=0
2(x-9)(x+12)=0
x-9=0
x=9 ans.
9+3=12 ans.
Proof:
9^2+12^2=15^2
81+144=225
225=225